The number of AstraZeneca Covid-19 vaccine recipients who had severe discomfort symptoms

found to be 15 per 1000 population. Assuming the number of people who received the

AstraZeneca Covid-19 vaccine had severe discomfort symptoms follows a Poisson distribution.


ii) Find the probability that in a randomly selected 200 population, there are 4 or 5 people

who received the AstraZeneca Covid-19 vaccine have severe discomfort symptoms.


i) Find the probability that in a randomly selected 15000 population, there are at least 181

people who received the AstraZeneca Covid-19 vaccine have severe discomfort

symptoms.


1
Expert's answer
2021-10-12T01:59:09-0400

ii)

"P(x=k)=\frac{\lambda^ke^{-\lambda}}{k!}"


"\lambda=" 15 per 1000 = 3 per 200


So, "\lambda=3"


"P(4\ or\ 5)=P(4)+P(5)=\frac{3^4e^{-3}}{4!}+\frac{3^5e^{-3}}{5!}=0.1680+0.1008=0.2688"


i)

"\lambda=15\cdot 15000/1000=225"


"P(X\le x)\le \frac{(e\lambda)^xe^{-\lambda}}{x^x}" for "x<\lambda"


"P(x\ge 181)=1-\frac{(225e)^{181}e^{-225}}{181^{181}}=0.0099"